= 2, and you We consider some function () whose domain of g is all real numbers except are given and g as follows: ) = et

Please help and show steps. Thank you.
Answer #1

Date: / Page no Given, q!(X) = ė (x-3) Co=23 (x) = x (x² - 6x +10) a (1-2) Partla) Either ex = 0 or (2-3) = 0 u=3 x-001 andDate:_, , Page no: At does (4-2)=0 not exists. the le = 2, 2 point where olla) does not exists. These are I called poles. -{9Date: / / Page no: - 2 At {x-2) = 0 x=2, 2, 2 These are the points where da) does not exists. Finding 9 (x) at [x=31 If gu(Date: 11 _Page no:_ • Point of Inflection of g(x) → These point can occur where the second derivative is zero. And we know th